# Symmetry re-breaking in an effective theory of quantum coarsening

### v1

@@ Abstract
We introduce a minimal model based on the Hamiltonian dynamics of classical spins to explain the experimental observations, which suggests that the phenomena are not characteristic of quantum systems. We further develop a mean-field approximation that allows for an analytical characterization of the observed oscillations. Our analysis reveals that these oscillations can already appear deep in the ordered phase, and are clearly distinct from the long-time coarsening dynamics. The onset of the latter is determined by the exponential growth of spatial fluctuations, which drive the system into an intermediate dynamical regime where long-range order is temporarily suppressed, before being re-established through coarsening. This gives the final state a sensitive dependence on quench parameters. This sequence of events unveils a robust phenomenon, which we term symmetry re-breaking, and should generically occur in the symmetry-broken phases of systems with Hamiltonian dynamics. We note that features of symmetry re-breaking have been previously observed in weakly perturbed fully connected Ising models [21, *Lerose2019Impact].
@@
We take the classical limit of this model by sending the spin representation from $S\!=\!1/2$ to $\infty$ while rescaling $\hat{S}_{i}^{\alpha}\to\hat{S}_{i}^{\alpha}/S$, thus obtaining the classical model [23]
@@
The commutation relations $\big{[}\hat{S}_{i}^{\alpha},\hat{S}_{j}^{\beta}\big{]}=i\hbar\delta_{ij}\epsilon^{\alpha\beta\gamma}\hat{S}_{i}^{\gamma}$ get mapped to the Poisson brackets $\big{\{}S_{i}^{\alpha},S_{j}^{\beta}\big{\}}=\delta_{ij}\epsilon^{\alpha\beta\gamma}S_{i}^{\gamma}$. Therefore, the classical system is endowed with a natural Hamiltonian dynamics $\partial_{t}\vec{S}=-(g\hat{x}+\hat{z}\sum_{j\in\partial i}S_{j}^{z})\wedge\vec{S}$, see Fig. 1(b), which conserves the total energy $H$. Owing to the bipartite nature of the lattice, this dynamics can be simulated to high accuracy with trotterized algorithms that conserve the energy to machine precision (see Supplement [24]).
The phase diagram of the classical model, Eq. (1), is shown in Fig. 1(a), as a function of the transverse field $g$ and the energy density $\varepsilon$ above the ground state. As in the quantum counterpart, a line of second-order phase transitions separates a paramagnetic (PM) phase at large $(g,\varepsilon)$ from a ferromagnetic (FM) one at small $(g,\varepsilon)$. The order parameter is the magnetization along $z$: $m=\frac{1}{L^{2}}\sum_{i}S^{z}_{i}$.
In the language of the Ising model, Eq. (1), the experimental observations [15] can be stated as follows. First, suppose the system is initialized in a domain-wall configuration with one macroscopic region in the “up” ground state ($m\!>\!0$) and the other in the “down” ground state ($m\!<\!0$). The interface between such two thermodynamically-large domains spans a length $O(L)$, so the total excess energy density vanishes in the thermodynamic limit as $O(1/L)$. Thus, the initial state is close to a ground state deep in the quantum FM phase. The experiment shows that the ensuing coarsening dynamics is curvature-driven: for a circular domain of radius $r$, the radius shrinks in time according to
@@
In the following, we address the above points, showing that they can be explained already within classical Hamiltonian dynamics. While the Ising model is different from the Rydberg Hamiltonian used in the experiments, the two can be treated with similar mean-field techniques and are expected to behave qualitatively alike.
Figure 1: Phase diagram and global dynamics in the classical transverse-field Ising model. (a) The Binder cumulant $B\!=\!3(1-\expectationvalue*{m^{4}}/3\expectationvalue*{m^{2}}^{2})/2$ distinguishes between the ordered ($B\!=\!1$) and disordered ($B\!=\!0$) phases. The red line shows the states accessed with the areal speed simulations. The black line shows the post-quench states accessed in the oscillation experiments which exhibit either ferromagnetic (solid), paramagnetic (dotted), or symmetry re-breaking (dashed) behaviors. The square dot marks the thermodynamic transition at $g=g_{c}$, and the round dot the dynamical transition at $g=g_{\mathrm{dyn}}$ in the mean-field model. The data was obtained via Monte Carlo simulations at a fixed temperature, for a $50\times 50$ lattice, and plotted as a function of the excess energy density $\varepsilon$. (b) The Bloch sphere of a classical spin $i$. Switching on a transverse field misaligns the net (exchange+applied) local field with the spin direction, leading to spin precession.
@@
To show that the classical limit $S\to\infty$ is able to capture features of the quantum spins-1/2, we investigate the (areal) speed of shrinking of a single minority domain, $v_{a}$, at different $(g,\varepsilon)$ in the FM phase. We take as initial state for the dynamics a circular domain of radius $r(t\!=\!0)\!=\!40$ lattice sites, as in Fig. 2(a). The transverse field $g$ is fixed through the Hamiltonian parameters. To fix $\varepsilon$, we first let the outside state be the “down” ground state ($S_{i}^{x}\!=\!g/4$, $S_{i}^{z}\!=\!-\sqrt{1-g^{2}/16}$), and the one inside the disk to be the “up” ($S_{i}^{x}\!=\!g/4$, $S_{i}^{z}\!=\!+\sqrt{1-g^{2}/16}$). We then slowly heat up the configuration by coupling it to an effective bath (via Monte Carlo sampling), until it reaches the target energy density above the ground state $\varepsilon$  [^2] The resulting configurations are not drawn from the Gibbs distribution as we do not run the Monte Carlo algorithm for long enough.. We finally time-evolve the resulting state: as shown in Fig. 2(b–c), the disk shrinks and eventually disappears.
Figure 2(d) shows that the area of the domain, measured by the number of spins with $S_{i}^{z}>0$ inside the original domain boundary, decreases linearly in time, consistent with curvature-driven coarsening: $dr^{2}/dt\simeq\mathrm{const.}=v_{a}$, see also Eq. (2). In Fig. 2(e), we also show that $v_{a}$ increases with $g$ within the FM phase: this is consistent with the experimental observations. However, closer to the critical line, one can see that $v_{a}$ decreases again as $g$ approaches $g_{c}$.
The two regimes—the areal speed increasing and then decreasing—reflect different physical mechanisms. Deep within the FM phase, $v_{a}$ increases with $g$ because $g$ defines the strength of the “kinetic term” $-g\sum_{i}S_{i}^{x}$: making it larger speeds up the spin dynamics generally, and in particular at the domain wall edges. In the quantum model, this heuristic can be made rigorous with prethermalization bounds as $g\to 0$ [27, 28, 29, 8, *Balducci2023Interface], but similar arguments can apply to classical models [30]. Separately, as $g$ approaches $g_{c}$, the vanishing of the domain-wall tension provides a mechanism for decreasing the areal speed $v_{a}$, as is evident in Fig. 2(e).
Figure 2: Areal speed of coarsening. An initial “up” circular domain, with the opposite magnetization as compared to its surroundings, is eroded as the system evolves with Hamiltonian dynamics [$g\!=\!1.5$ and times $t\!=\!0$,200,400 for panels (a),(b),(c) respectively]. (d–e) The speed $v_{a}$ at which this happens increases with $g$ starting from $g\!=\!0$, until it decreases in the proximity of the phase transition. The excess energy density is fixed to $\varepsilon\!=\!0.2$, as marked in Fig. 1.
@@
We now study quench dynamics within the ordered phase. To compare with experiments, we prepare the system at $g\!=\!0$ and fixed excess energy density $\varepsilon_{0}\gtrsim 0$ by “warming up” gradually the perfect ferromagnet $\vec{S}_{i}\!=\!+\hat{z}$ via Monte Carlo sampling. A nonzero energy density is essential, as the classical model’s ground state is uniform and, in the absence of fluctuations, the uniform (zero-momentum) mode decouples from the dynamics. We perform a sudden quench by evolving the system with the Hamiltonian $H[g]$ for $g\!>\!0$. The locus of the excess energy density of the post-quench states is marked with the black line in Fig. 1 for the case $\varepsilon_{0}\!=\!0$; the lines corresponding to $\varepsilon_{0}\!\gtrsim\!0$ lie above the displayed one.
Figure 3 shows that the quench protocol induces large-scale oscillations of the order parameter $m$, at $\varepsilon_{0}=0.2$. Deep in the FM phase, these oscillations remain confined to $m\!>\!0$, indicating the persistence of the initial symmetry-broken state (Fig. 3(a)), while $m$ takes both signs in the PM, eventually relaxing to zero (Fig. 3(d)). However, there is a finite region within the FM in which $m$ changes sign, either once (Fig. 3(b)) or several times (Fig. 3(c)) before relaxing to a value which slowly drifts with time towards the equilibrium value (Fig. 4(a)).
Figure 3: Post-quench oscillations. (a–d) After the quench, the magnetization (colored lines) oscillates at the frequency set by the mean-field solution (pale grey lines), while getting damped. The plots are respectively for $g=1.8,2.0,2.2,2.5$; the thermodynamic phase transition is at $g\simeq 2.4$. (e) Comparison of the oscillation frequencies with the mean-field prediction in gray, showing excellent agreement. The frequency minimum does not correspond to the Ising transition (vertical dotted line). Notice the absence of damping $\gamma$ in the mean field theory. (f) Bloch sphere representation of each spin in the uniform state, showing a trajectory confined to one half of the Bloch sphere, one crossing to the other half of the sphere, and the separatrix between the two.
These features can be understood in a mean-field (MF) analysis of the post-quench dynamics. The first step is to notice that, when a spatially uniform configuration $\vec{S}_{i}\equiv\vec{S}$ is evolved with the classical Hamiltonian, the zero-momentum mode decouples completely. The effective equations of motion describing its dynamics are, self-consistently, of MF type:
@@
where the factor of 4 reflects the coordination number of the square lattice. These equations of motion correspond to those of the fully-connected, Lipkin-Meshkov-Glick (LMG) Hamiltonian $H=-4(S^{z})^{2}-gS^{x}$ [31]. Since the phase space of the MF model is two-dimensional (i.e. the surface of the unit sphere) and there is one conserved quantity (the energy), the dynamics is regular and the spin precesses along closed loops, see Fig. 3(f). By solving Eq. (3), one can extract the frequency $\omega(g)$ of the oscillations, which agrees quantitatively with the frequency extracted from the oscillations in the 2D model, see Fig. 3(e). Observe that the minimum of the oscillation frequency at $g\!=\!2$ does not occur at the thermodynamic phase transition of the 2D model ($g_{c}\simeq 2.4$ for the energy density used). Instead, a dynamical phase transition occurs at $g\!=\!g_{\mathrm{dyn}}\!=\!2$ in the LMG model [32, 21, *Lerose2019Impact], manifesting in the Bloch sphere as a separatrix (Fig. 3(f)). To one side of the separatrix ($g\!<\!2$), the magnetization is confined to the upper-half sphere $S^{z}\!>\!0$, while on the other side ($g\!>\!2$) $S^{z}$ takes both signs. This is also reflected in the maximum of the phenomenological damping rate $\gamma$—as obtained by fitting the underdamped oscillations—being at $g\!=\!2$, see Fig. 3(e).
@@
We are now in position to explain quenches in the FM phase, in which oscillations drive the magnetization to near zero, or even reverse its sign (Figs. 3(b–c)) at late times. Observe that in the LMG model, the quench to $g\!=\!2$ (Fig. 3(b)) corresponds to a separatrix trajectory that terminates at the unstable PM fixed point $\vec{S}=\hat{x}$. In practice, the LMG description holds locally for the 2D model, and small fluctuations shift the local MF to either one or the other side of the separatrix. The result is that some regions of the 2D lattice end up “bouncing back” to $S_{i}^{z}>0$, while others cross over to $S_{i}^{z}<0$. As the order set by the initial condition is lost, self-consistently the MF description ceases to be valid. As the system is in the FM phase in equilibrium, it coarsens again without ever leaving the ordered phase, to re-establish the magnitude of the magnetization in thermodynamic equilibrium. We term this process symmetry re-breaking.
Symmetry re-breaking takes place in the whole region $g_{\mathrm{dyn}}\!\leq\!g\!<\!g_{c}$. For $g\!>\!g_{\mathrm{dyn}}\!=\!2$, the magnetization performs a few oscillations. The oscillations feature a non-exponential relaxation, which terminates consistently at the same time across different realizations of the initial (noisy) state, and at a non-thermal value for $m$. We investigate this feature in detail in the next section. Figure 4(a) shows that, at the end of the oscillations, the system is partitioned into different domains of “up” and “down” spins, the relative abundance of which seems fixed to a certain ratio. These domains represent the seed for coarsening dynamics, which drives the formation of larger and larger domains. Due to the imbalance of the number of “up” and “down” spins in the post-oscillation states, the system ends with either a positive or negative magnetization $m_{\infty}$ in the infinite-time limit, as shown in Fig. 4(b–c). The dependence of $m_{\infty}$ on the quench parameters $g,\varepsilon$ seems highly complicated.
Figure 4: Symmetry re-breaking. (a) After the magnetization oscillations die out, the snapshot shows that the system is locally ordered, but globally broken into domains (colorbar identical to Fig. 2). These domains coarsen in time, eventually leading to the reestablishment of a global magnetization at one of its two equilibrium values. (b–c) The sign of the steady-state magnetization $m_{\infty}$ is determined by the parity of the number of oscillations the system undergoes before they are damped out. This number increases with decreasing energy density of the pre-quench state.
@@
Symmetry re-breaking is a mechanism driven by an interplay of a mean field (MF) and spatially-varying fluctuations on top of it. The behavior of the MF is fully captured by Eq. (3); here we analyze the fluctuations by developing a self-consistent gaussian theory thereof. Doing so for the $S=\infty$ Ising model results in a complicated set of equations, that are difficult to analyze [22]. For this reason, we investigate the same phenomenon in a simpler $\phi^{4}$ theory in 2D. As a side result, this showcases the generality of our findings beyond the specific choice of model.
@@
where $\nabla$ represents the lattice derivative: $(\nabla\phi)_{(x,y)}\equiv(\phi_{(x+1,y)}-\phi_{(x,y)},\phi_{(x,y+1)}-\phi_{(x,y)})$ for $i\!=\!(x,y)$, $\phi_{i}$ is a scalar field, and $\pi_{i}$ the conjugate momentum. The presence of the conjugate momenta $\pi_{i}$ ensures that the dynamics is Hamiltonian. As described in the Supplement [24], we single out the MF by passing to Fourier space $\phi_{i}\to\tilde{\phi}_{k}$: the MF corresponds to the $k\!=\!0$ mode, while the $k\neq 0$ modes encode spatial variations in the scalar field. Then, we approximate the $k\neq 0$ modes as gaussian fluctuations on top of the MF by replacing $\tilde{\phi}^{4}\to 3\expectationvalue*{\tilde{\phi}^{2}}\tilde{\phi}^{2}$ using Wick’s theorem. After these approximations, the equations of motion become
@@
We trace the exponential growth of fluctuations to the presence of an “inverted mass” term. Namely, Eq. (5b) for low momenta acquires the form $\partial_{t}^{2}\tilde{\phi}_{k}\approx-(r+3u\,\tilde{\phi}_{0}^{2})\tilde{\phi}_{k}$, and fluctuations grow exponentially in time since $\tilde{\phi}_{0}^{2}$ is driven close to 0 and $r\!<\!0$. The exponential growth of fluctuations explains why the observed decay of the MF mode is not exponential: the oscillations are rapidly cut off when the fluctuations become of order the value of mean field, but not before. One can finally notice that the full interacting model and the gaussian approximation behave differently after the oscillations have decayed. In fact, the 2D interacting model undergoes energy-conserving coarsening dynamics, as shown for the classical Ising model, while the gaussian model remains non-thermal up to infinite time [33, 34].
Figure 5: Role of fluctuations. Order parameter oscillations (a) and growth of fluctuations (b) for the 2D $\phi^{4}$ theory, Eq. (4), for a quench $r_{i}\!=\!-3.9$ to $r\!=\!-2$. The exact dynamics (solid blue lines) agrees with the gaussian decoupling (Eq. (5), pale gray lines) until the fluctuations become of the same order of the mean field. Panels (c) and (d) similarly show a quench into the symmetry re-breaking region ($r_{i}\!=\!-4.1$, $r\!=\!-2$), for a case where the asymptotic value of the order parameter has the opposite sign.
@@
We have provided a transparent explanation of two notable experimental observations in a quantum coarsening experiment [15]. First, we have shown that the observed speed-up is explained by the tuning parameter playing a dual role, i.e. tuning the system closer to the phase transition but also tuning the strength of the kinetic term. Second, we have thoroughly characterized the observed order parameter oscillations, predicting their shape and persistence. In doing so we have unveiled the generic phenomenon of symmetry re-breaking, in which quenches entirely within an ordered phase can lead to the destruction of order at intermediate times, and its reemergence with either asymptotic sign of the order parameter at long times. Our results reveal that, even in dimension as low as $d\!=\!2$, and for the “most quantum” spins-1/2, semiclassical analyses are a viable tool for disentangling purely quantum effects from other many-body phenomena. Further studies are clearly called for, on any of the existing platforms, in particular to investigate symmetry re-breaking, and the detailed dynamics in the critical region.

### v2

@@ Abstract
We introduce a minimal model based on the Hamiltonian dynamics of classical spins to explain the experimental observations, which suggests that the phenomena are not characteristic of quantum systems. We further develop a mean-field (MF) approximation that allows for an analytical characterization of the observed oscillations. Our analysis reveals that these oscillations can already appear deep in the ordered phase, and are clearly distinct from the long-time coarsening dynamics. The onset of the latter is determined by the exponential growth of spatial fluctuations, which drive the system into an intermediate dynamical regime where long-range order is temporarily suppressed, before being re-established through coarsening. This gives the final state a sensitive dependence on quench parameters. This sequence of events unveils a robust phenomenon, which we term symmetry re-breaking, and should generically occur in the symmetry-broken phases of systems with Hamiltonian dynamics. We note that features of symmetry re-breaking have been previously observed in weakly perturbed fully connected Ising models [21, *Lerose2019Impact].
@@
We take the classical limit of this model by sending the spin representation from S=1/2S\!=\!1/2 to $\infty$ while rescaling $\hat{S}_{i}^{\alpha}\to\hat{S}_{i}^{\alpha}/S$, thus obtaining the classical model [23]
@@
The commutation relations $\big[\hat{S}_{i}^{\alpha},\hat{S}_{j}^{\beta}\big]=i\hbar\delta_{ij}\epsilon^{\alpha\beta\gamma}\hat{S}_{i}^{\gamma}$ get mapped to the Poisson brackets $\big\{S_{i}^{\alpha},S_{j}^{\beta}\big\}=\delta_{ij}\epsilon^{\alpha\beta\gamma}S_{i}^{\gamma}$. Therefore, the classical system is endowed with a natural Hamiltonian dynamics $\partial_{t}\vec{S}_{i}=-(g\hat{x}+\hat{z}\sum_{j\in\partial i}S_{j}^{z})\wedge\vec{S}_{i}$, see Fig. 1(b), which conserves the total energy $H$. Owing to the bipartite nature of the lattice, this dynamics can be simulated to high accuracy with trotterized algorithms that conserve the energy to machine precision (see End Matter). The initial conditions of the dynamics are specified case by case.
The phase diagram of the classical model, Eq. (1), is shown in Fig. 1(a), as a function of the transverse field $g$ and the excess energy density $\varepsilon$, which is the energy density in excess of that in the ground state. As in the quantum counterpart, a line of second-order phase transitions separates a paramagnetic (PM) phase at large $(g,\varepsilon)$ from a ferromagnetic (FM) one at small $(g,\varepsilon)$. The order parameter is the magnetization along $z$: $m=\frac{1}{L^{2}}\sum_{i}S^{z}_{i}$.
In the language of the Ising model, Eq. (1), the experimental observations [15] can be stated as follows. First, suppose the system is initialized in a domain-wall configuration with one macroscopic region in the “up” ground state (m>0m\!>\!0) and the other in the “down” ground state (m<0m\!<\!0). The interface between such two thermodynamically-large domains spans a length $O(L)$, so the total excess energy density vanishes in the thermodynamic limit as $O(1/L)$. Thus, the initial state is close to a ground state deep in the quantum FM phase. The experiment shows that the ensuing coarsening dynamics is curvature-driven: for a circular domain of radius $r$, the radius shrinks in time according to
@@
In the following, we address the above points, showing that they can be explained already within classical Hamiltonian dynamics. While the Ising model is different from the Rydberg Hamiltonian used in the experiments, the two can be treated with similar MF techniques and are expected to behave qualitatively alike.
Figure 1: Phase diagram and global dynamics in the classical transverse-field Ising model. (a) The Binder cumulant $B\!=\!3(1-\expectationvalue*{m^{4}}/3\expectationvalue*{m^{2}}^{2})/2$ distinguishes between the ordered (B=1B\!=\!1) and disordered (B=0B\!=\!0) phases. The red line shows the states accessed with the areal speed simulations. The black line shows the post-quench states accessed in the oscillation experiments which exhibit either ferromagnetic (solid), paramagnetic (dotted), or symmetry re-breaking (dashed) behaviors. The square dot marks the thermodynamic transition at $g=g_{c}$, and the round dot the dynamical transition at $g=g_{\mathrm{dyn}}$ in the mean-field model. The data was obtained via Monte Carlo simulations at a fixed temperature, for a $\times50$ lattice, and plotted as a function of the excess energy density $\varepsilon$. (b) The Bloch sphere of a classical spin ii. Switching on a transverse field misaligns the net (exchange+applied) local field with the spin direction, leading to spin precession.
@@
To show that the classical limit $S\to\infty$ is able to capture features of the quantum spins-1/2, we investigate the (areal) speed of shrinking of a single minority domain, $v_{a}$, for different $(g,\varepsilon)$ in the FM phase. We take as initial state for the dynamics a circular domain, as shown in Fig. 2(a)—a more detailed description is in the End Matter. During time evolution the disk shrinks and eventually disappears [Fig. 2(b–c)].
Figure 2(d) shows that the area of the domain, measured by the number of spins with $S_{i}^{z}>0$ inside the original domain boundary, decreases linearly in time, consistent with curvature-driven coarsening: $dr^{2}/dt\simeq\mathrm{const.}=v_{a}$, see also Eq. (2). In Fig. 2(e), we also show that $v_{a}$ increases with $g$ already well within the FM phase: this is consistent with the experimental observations. However, closer to the critical line, one can see that $v_{a}$ decreases again as $g$ approaches $g_{c}$.
The two regimes—the areal speed increasing and then decreasing—reflect different physical mechanisms. Deep within the FM phase, $v_{a}$ increases with $g$ because $g$ defines the strength of the “kinetic term” $g\sum_{i}S_{i}^{x}$: making it larger speeds up the spin dynamics generally, and in particular at the domain wall edges. In the quantum model, this heuristic can be made rigorous with prethermalization bounds as $g\to 0$ [25, 26, 27, 8, *Balducci2023Interface], but similar arguments can apply to classical models [28]. Separately, as $g$ approaches $g_{c}$, the vanishing of the domain-wall tension provides a mechanism for decreasing the areal speed $v_{a}$, as is evident in Fig. 2(e).
Figure 2: Areal speed of coarsening. An initial “up” circular domain, with the opposite magnetization as compared to its surroundings, is eroded as the system evolves with Hamiltonian dynamics [g=2g\!=\!2 and times t=0t\!=\!0,250,500 for panels (a),(b),(c) respectively]. (d–e) The speed $v_{a}$ at which this happens increases with $g$ starting from g=0g\!=\!0, until it decreases in the proximity of the phase transition. The excess energy density is fixed to $\varepsilon\!=\!0.2$, as marked in Fig. 1. A lattice of size $\times400$ is used for the simulations.
@@
We now study quench dynamics within the ordered phase. To compare with experiments, we prepare the system at g=0g\!=\!0 and fixed excess energy density $\varepsilon_{0}\gtrsim 0$. A nonzero excess energy density is essential, as the classical model’s ground state is uniform and, in the absence of fluctuations, the uniform (zero-momentum) mode decouples from the dynamics. We perform a sudden quench by evolving the system with the Hamiltonian $H[g]$ for g>0g\!>\!0. The locus of the excess energy density of the post-quench states is marked with the black line in Fig. 1 for the case $\varepsilon_{0}\!=\!0$; the lines corresponding to $\varepsilon_{0}\!\gtrsim\!0$ lie above the displayed one.
Figure 3 shows that the quench protocol induces large-scale oscillations of the order parameter $m$, at $\varepsilon_{0}=0.2$. Deep in the FM phase, these oscillations remain confined to m>0m\!>\!0, indicating the persistence of the initial symmetry-broken state (Fig. 3(a)), while $m$ takes both signs in the PM, eventually relaxing to zero (Fig. 3(d)). However, there is a finite region within the FM in which $m$ changes sign, either once (Fig. 3(b)) or several times (Fig. 3(c)) before relaxing to a value which slowly drifts with time towards the equilibrium value (Fig. 4(a)).
Figure 3: Post-quench oscillations. (a–d) After the quench, the magnetization (colored lines) oscillates at the frequency set by the mean-field solution (pale grey lines), while getting damped. The plots are respectively for $g=1.8,2.0,2.2,2.5$; the thermodynamic phase transition is at $g_{c}\simeq 2.4$. (e) Comparison of the oscillation frequencies with the mean-field prediction in gray, showing excellent agreement. The frequency minimum does not correspond to the Ising transition (vertical dotted line). Notice the absence of damping $\gamma$ in the mean field theory. (f) Bloch sphere representation of each spin in the uniform state, showing a trajectory confined to one half of the Bloch sphere, one crossing to the other half of the sphere, and the separatrix between the two. A lattice of size $\times200$ is used for the simulations.
These features can be understood in a MF analysis of the post-quench dynamics. The first step is to notice that, when a spatially uniform configuration $\vec{S}_{i}\equiv\vec{S}$ is evolved with the classical Hamiltonian, the zero-momentum mode decouples completely. The effective equations of motion describing its dynamics are, self-consistently, of MF type:
@@
where the factor of 4 reflects the coordination number of the square lattice. These equations of motion correspond to those of the fully-connected, Lipkin-Meshkov-Glick (LMG) Hamiltonian $H=-2(S^{z})^{2}-gS^{x}$ [29]. Since the phase space of the MF model is two-dimensional (i.e. the surface of the unit sphere) and there is one conserved quantity (the energy), the dynamics is regular and the spin precesses along closed loops, see Fig. 3(f). The LMG oscillation frequency $\omega(g)$, computed in the End Matter, agrees quantitatively with the frequency extracted from the oscillations in the 2D model, see Fig. 3(e). Observe that the minimum of the oscillation frequency at g=2g\!=\!2 does not occur at the thermodynamic phase transition of the 2D model ($g_{c}\simeq 2.4$ for the energy density used). Instead, a dynamical phase transition occurs at $g\!=\!g_{\mathrm{dyn}}\!=\!2$ in the LMG model [30, 21, *Lerose2019Impact], manifesting in the Bloch sphere as a separatrix (Fig. 3(f)). To one side of the separatrix (g<2g\!<\!2), the magnetization is confined to the upper-half sphere $S^{z}\!>\!0$, while on the other side (g>2g\!>\!2) $S^{z}$ takes both signs. This is also reflected in the maximum of the phenomenological damping rate $\gamma$—as obtained by fitting the underdamped oscillations—being at g=2g\!=\!2, see Fig. 3(e).
@@
We are now in position to explain quenches in the FM phase, in which oscillations drive the magnetization to near zero, or even reverse its sign (Figs. 3(b–c)) at late times. Observe that in the LMG model, the quench to g=2g\!=\!2 (Fig. 3(b)) corresponds to a separatrix trajectory that terminates at the unstable PM fixed point $\vec{S}=\hat{x}$. In practice, the LMG description holds locally for the 2D model, and small fluctuations shift the local MF to either one or the other side of the separatrix. The result is that some regions of the 2D lattice end up “bouncing back” to $S_{i}^{z}>0$, while others cross over to $S_{i}^{z}<0$. As the order set by the initial condition is lost, self-consistently the MF description ceases to be valid. As the system is in the FM phase in equilibrium, it coarsens again without ever leaving the ordered phase, to re-establish the magnitude of the magnetization in thermodynamic equilibrium. We term this process symmetry re-breaking.
Symmetry re-breaking takes place in the whole region $g_{\mathrm{dyn}}\!\leq\!g\!<\!g_{c}$. For $g\!>\!g_{\mathrm{dyn}}\!=\!2$, the magnetization performs a few oscillations. The oscillations feature a non-exponential relaxation, which terminates consistently at the same time across different realizations of the initial (noisy) state, and at a non-thermal value for $m$. We investigate this feature in detail in the next section. Figure 4(a) shows that, at the end of the oscillations, the system is partitioned into different domains of “up” and “down” spins, the relative abundance of which seems fixed to a certain ratio. These domains represent the seed for coarsening dynamics, which drives the formation of larger and larger domains. Due to the imbalance of the number of “up” and “down” spins in the post-oscillation states, the system ends with either a positive or negative magnetization $m_{\infty}$ in the infinite-time limit, as shown in Fig. 4(b–c). The dependence of the sign of $m_{\infty}$ on the quench parameters $g,\varepsilon$ seems highly complicated, while its absolute value is set by the microcanonical ensemble.
Figure 4: Symmetry re-breaking. (a) After the magnetization oscillations die out, the snapshot shows that the system is locally ordered, but globally broken into domains (colorbar identical to Fig. 2). These domains coarsen in time, eventually leading to the reestablishment of a global magnetization at one of its two equilibrium values. (b–c) The sign of the steady-state magnetization $m_{\infty}$ is determined by the parity of the number of oscillations the system undergoes before they are damped out. This number increases with decreasing energy density of the pre-quench state. A lattice of size $\times200$ is used for the simulations.
@@
Symmetry re-breaking is a mechanism driven by an interplay of a MF and spatially-varying fluctuations on top of it. The behavior of the MF is fully captured by Eq. (3); here we analyze the fluctuations by developing a self-consistent gaussian theory thereof. Doing so for the $S=\infty$ Ising model results in a complicated set of equations, that are difficult to analyze [22]. For this reason, we investigate the same phenomenon in a simpler $\phi^{4}$ theory in 2D. As a side result, this showcases the generality of our findings beyond the specific choice of model.
@@
where $\nabla$ represents the lattice derivative: $(\nabla\phi)_{(x,y)}\equiv(\phi_{(x+1,y)}-\phi_{(x,y)},\phi_{(x,y+1)}-\phi_{(x,y)})$ for i=(x,y)i\!=\!(x,y), $\phi_{i}$ is a scalar field, and $\pi_{i}$ the conjugate momentum. The presence of the conjugate momenta $\pi_{i}$ ensures that the dynamics is Hamiltonian. As described in the End Matter, we single out the MF by passing to Fourier space $\phi_{i}\to\tilde{\phi}_{k}$: the MF corresponds to the k=0k\!=\!0 mode, while the $k\neq 0$ modes encode spatial variations in the scalar field. Then, we approximate the $k\neq 0$ modes as gaussian fluctuations on top of the MF by replacing $\tilde{\phi}^{4}\to 3\expectationvalue*{\tilde{\phi}^{2}}\tilde{\phi}^{2}$ using Wick’s theorem. After these approximations, the equations of motion become
@@
We trace the exponential growth of fluctuations to the presence of an “inverted mass” term. Namely, Eq. (5b) for low momenta acquires the form $\partial_{t}^{2}\tilde{\phi}_{k}\approx-(r+3u\,\tilde{\phi}_{0}^{2})\tilde{\phi}_{k}$: since $\tilde{\phi}_{0}^{2}$ passes through 0 and r<0r\!<\!0, the equation takes periodically the form of an inverted oscillator, making fluctuations grow exponentially in time. The exponential growth of fluctuations explains why the observed decay of the MF mode is not exponential: the oscillations are rapidly cut off when the fluctuations become of the same order as the value of the MF, but not before. One can finally notice that the full interacting model and the gaussian approximation behave differently after the oscillations have decayed. In fact, the 2D interacting model undergoes energy-conserving coarsening dynamics, as shown for the classical Ising model, while the gaussian model remains non-thermal up to infinite time [31, 32, 33, 34].
Figure 5: Role of fluctuations. Order parameter oscillations (a) and growth of fluctuations (b) for the 2D $\phi^{4}$ theory, Eq. (4), for a quench $r_{i}\!=\!-3.9$ to r=2r\!=\!-2. The exact dynamics (solid blue lines) agrees with the gaussian decoupling (Eq. (5), pale gray lines) until the fluctuations become of the same order of the mean field. Panels (c) and (d) similarly show a quench into the symmetry re-breaking region ($r_{i}\!=\!-4.1$, r=2r\!=\!-2), for a case where the asymptotic value of the order parameter has the opposite sign. A lattice of size $\times201$ is used for the simulations.
@@
We have provided a transparent explanation of two notable experimental observations in a quantum coarsening experiment [15]. First, we have argued that the observed speed-up is explained by the tuning parameter playing a dual role, i.e. tuning the system closer to the phase transition but also tuning the strength of the kinetic term. While our argument was heuristic, it can be made more rigorous through a Schrieffer-Wolff-like transformation. Second, we have thoroughly characterized the observed order parameter oscillations, predicting their shape and persistence. In doing so we have unveiled the generic phenomenon of symmetry re-breaking, in which quenches entirely within an ordered phase can lead to the destruction of order at intermediate times, and its reemergence with either sign at asymptotically long times. Our results reveal that, even in dimension as low as d=2d\!=\!2, and for the “most quantum” spins-1/2, semiclassical analyses are a viable tool for disentangling purely quantum effects from other many-body phenomena. Further studies are clearly called for, on any of the existing platforms, in particular to investigate symmetry re-breaking, and the detailed dynamics in the critical region. A first, viable analysis is the use of a $1/S$ expansion around the classical trajectories studied in this work.